# Convergence to equilibrium for a degenerate three species reaction-diffusion system

### v1

@@ 2. The case of $d_{b}=0$
$\leq$ $\|a_{0}+c_{0}\|_{L^{p}(\Omega)}\|\phi(0)\|_{L^{q}(\Omega)}+\|a\|_{L^{p}(\Omega_{0,2})}\|\partial_{t}\phi\|_{L^{q}(\Omega_{0,2})}+d_{a}\|a\|_{L^{p}(\Omega_{0,2})}\|\Delta\phi\|_{L^{q}(\Omega_{0,2})}.$
Choose $q=\frac{4.5}{4.5-1}$,i.e., Hölder conjugate of $4.5$,duality estimate yields
$\|c\|_{L^{4.5}(\Omega_{0,2})}\leq\|a_{0}+c_{0}\|_{L^{4.5}(\Omega)}+2C_{q,d_{c}}\|a\|_{L^{4.5}(\Omega_{0,2})}.$
Choose $K_{0,1}=\max\left\{(C_{IE}+2C_{q,d_{c}})(\Omega,d_{a})K_{7}+\|a_{0}+c_{0}\|_{L^{4.5}(\Omega)}\right\}$, we have the following estimate
$\|a\|_{L^{4.5}(\Omega_{0,1})}+\|c\|_{L^{4.5}(\Omega_{0,1})}\leq K_{0,1}.$
@@
$\partial_{t}\left(\phi_{\tau}a\right)-d_{a}\Delta\left(\phi_{\tau}a\right)&=a\partial_{t}\phi_{\tau}+\phi_{\tau}(c-ab)\qquad\mbox{ in}\Omega_{\tau,\tau+2}\\ \phi_{\tau}\nabla a\cdot n&=0\qquad\qquad\qquad\qquad\qquad\ \mbox{ on }\partial\Omega_{\tau,\tau+2}\\ \phi_{\tau}(\tau)a(\tau,x)&=0\qquad\qquad\qquad\qquad\qquad\ \ \ \mbox{ in }\Omega\end{aligned}\right.$
@@
[…]
@@
Employing the integrability estimate for $\zeta$ available from [CDF14, Lemma 3.3] (see Theorem 11 from the Appendix for the precise statement), we deduce that
@@
$\partial_{t}\left(\phi_{\tau}c\right)-d_{c}\Delta\left(\phi_{\tau}c\right)=&c\partial_{t}\phi_{\tau}+\phi_{\tau}(ab-c)\qquad\mbox{ in}\Omega_{\tau,\tau+2}\\ \phi_{\tau}\nabla c\cdot n&=0\qquad\qquad\qquad\qquad\ \mbox{ on }\partial\Omega_{\tau,\tau+2}\\ \phi_{\tau}(\tau)c(\tau,x)&=0\qquad\qquad\qquad\qquad\ \ \ \mbox{ in }\Omega\end{aligned}\right.$
where $\phi_{\tau}(s):=\phi(s-\tau)$ for $s\in[\tau,\infty)$ with $\tau>0$. Thanks to the positivity of $c$, employing the maximum principle for the heat equation and the integrability estimate from [CDF14, Lemma 3.3] (see Theorem 11 from the Appendix for the precise statement), we arrive at
@@
$\left\|b(t,\cdot)\right\|^{\frac{9}{4}}_{\mathrm{L}^{\frac{9}{4}}(\Omega)}$ $\leq 2^{\frac{5}{4}}(1+t)^{\frac{5}{4}}\left(\left\|b_{0}\right\|_{\mathrm{L}^{\infty}(\Omega)}^{\frac{9}{4}}\,\left|\Omega\right|+(1+t)^{\frac{47}{20}}K_{12}^{\frac{9}{4}}\right)\leq K_{13}(1+t)^{\frac{18}{5}}$
@@
where we have used the bounds obtained earlier. As $\frac{7}{2}<\frac{(5)\frac{3}{2}}{5-3}$, employing the maximum principle for the heat equation and the integrability estimate from [CDF14, Lemma 3.3] (see Theorem 11 from the Appendix for the precise statement) as before, we get
@@
In our above proof Gagliardo-Nirenberg inequality played an important role. We will use the same technique to get this   $L^{\frac{3}{2}}(\Omega)$ integral estimation of  $b$  for N=1,2 case.
@@
Let $a,.b,c$ be the solution of the degenerate system-(2) in dimension $N=1,2$, then there exists constants $\hat{K},K_{c}>0$ and some $\mu_{c}\in\mathbb{N}$, independent of time, such that:
@@
Similar as in the proof of proposition-3, we are going to use Gagliardo-Nirenberg inequality which says there exists a constant depending on the domain $C_{2}=C_{2}(\Omega,N)$ N be the dimension , for $1\leq p,q,r\leq+\infty$
$\|D^{j}\sqrt{a}\|_{L^{p}(\Omega)}\leq C_{2}\|D^{m}\sqrt{a}\|_{L^{r}(\Omega)}^{\alpha}\|\sqrt{a}\|_{L^{q}(\Omega)}^{1-\alpha}+\tilde{C_{2}}(\Omega,N,r,q)\|\sqrt{a}\|_{L^{q}(\Omega)}$
where,   $\frac{1}{p}=\frac{j}{N}+\big{(}\frac{1}{r}-\frac{m}{N}\big{)}\alpha\ +\frac{1-\alpha}{q}$   and   $\frac{j}{N}\leq\alpha\leq 1$
For N=2 take  $r=2,m=1,j=0,q=2$ and $\alpha=\frac{1}{2}$ we get
$\|\sqrt{a}\|_{L^{4}(\Omega)}^{4}\leq 2^{3}C_{2}^{4}\|\nabla\sqrt{a}\|_{L^{2}(\Omega)}^{2}\|\sqrt{a}\|_{L^{2}(\Omega)}^{2}+2^{3}\tilde{C_{2}}^{4}\|\sqrt{a}\|_{L^{2}(\Omega)}^{4}$
We have just like in theorem 3   $\forall\tau\geq 0$
$\int_{\tau}^{\tau+2}\|\sqrt{a}\|_{W^{1,2}(\Omega)}^{2}\leq K_{1};\ \ \int_{\tau}^{\tau+2}\|\sqrt{c}\|_{W^{1,2}(\Omega)}^{2}\leq K_{1}$
Where $K_{1}=\frac{2E(a_{0},b_{0},c_{0})}{d_{a}}+\frac{2E(a_{0},b_{0},c_{0})}{d_{c}}+2M_{1}\Omega+2M_{2}\Omega$. Integrating from  $\tau$  to  $\tau+2$
$\int_{\tau}^{\tau+2}\|\sqrt{a}\|_{L^{4}(\Omega)}^{4}\leq 2^{4}C_{2}^{4}M_{1}^{2}|\Omega|^{2}K_{1}+2^{4}\tilde{C_{2}}^{4}M_{1}^{4};$
$\int_{\tau}^{\tau+2}\|\sqrt{c}\|_{L^{4}(\Omega)}^{4}\leq 2^{4}C_{2}^{4}M_{1}^{2}|\Omega|^{2}K_{1}+2^{4}\tilde{C_{2}}^{4}M_{1}^{4}$
@@
Take  $K_{2}=(2^{4}C_{2}^{4}M_{1}^{2}|\Omega|^{2}K_{1}+2^{4}\tilde{C_{2}}^{4}M_{1}^{4})^{\frac{1}{2}}$, we have, $\forall\tau\geq 0$
$\|a\|_{L^{2}(\Omega_{\tau,\tau+2})}\leq K_{2},\qquad\|c\|_{L^{2}(\Omega_{\tau,\tau+2})}\leq K_{2}$
Applying Hölder inequality we get for all  $\tau\geq 0$,
(30) $\|a\|_{L^{\frac{5}{3}}(\Omega_{\tau,\tau+2})}\leq 2K_{2}|\Omega|^{\frac{1}{6}},\qquad\|c\|_{L^{\frac{5}{3}}(\Omega_{\tau,\tau+2})}\leq 2K_{2}|\Omega|^{\frac{1}{6}}$
Rest of the proof is just like the proof of proposition 3. We will get  $\hat{K}>0$ such that for  $t\geq 0$
$\|b\|_{L^{\frac{3}{2}}(\Omega)}\leq\hat{K}(1+t)^{\frac{5}{6}}$
Now for N=1 we take, $r=2,m=1,j=0,q=2$ and $\alpha=\frac{1}{3}$ and like in the previous N=2 case a similar approach gives us $\exists\ \hat{K}>0,K_{c}>0$ and $\mu_{c}\in\mathbb{N}$, such that for $\tau\geq 0$
$\|b\|_{L^{\frac{3}{2}}(\Omega)}$ $\leq\hat{K}(1+\tau)^{\frac{5}{6}};\qquad\ \tau\geq 0$
$\|a\|_{L^{3.5}(\Omega_{\tau,\tau+1})}$ $\leq K_{c}(1+\tau)^{\mu_{c}};\qquad\tau\geq 0$
$\|c\|_{L^{3.5}(\Omega_{\tau,\tau+1})}$ $\leq K_{c}(1+\tau)^{\mu_{c}};\qquad\tau\geq 0$
@@
Now above proposition-1 and proposition-3 will help us to relate entropy dissipation with the missing term  $\delta_{B}$. Furthermore the integral estimates of $c$ will provide us polynomial growth for the solutions too. We begin by the polynomial growth in the next theorem.
@@
Let $a,b,c$ be the solution of degenerate system-(2). Furthermore let for $p>\frac{N+2}{2}$, there exists a constant $K_{c}>0$, and some $\mu_{c}\in\mathbb{N}$, independent of time, such that:
$\|a\|_{L^{p}(\Omega_{\tau,\tau+1})}+\|c\|_{L^{p}(\Omega_{\tau,\tau+1})}\leq K_{c}(1+\tau)^{\mu_{c}},\quad\forall\tau\geq 0$
Then there exists constant $K_{\infty}$ and $\mu\in\mathbb{N}$, such that:
@@
Cnsidering $\phi:[0,\infty)\rightarrow[0,1]$ be a smooth function such that $\phi(0)=0$; $\phi\big{|}_{[1,\infty)}=1\text{\ and \ }\phi^{\prime}\in[0,M_{\phi}]$.
Let’s denote $\phi_{\tau}(s)=\phi(s-\tau)$, then:
@@
$\partial_{t}\phi_{\tau}a-d_{a}\Delta\phi_{\tau}a=&\phi^{\prime}_{\tau}a+\phi_{\tau}(c-ab)\qquad\mbox{ in }\Omega_{\tau,\tau+2}\\ \nabla\phi_{\tau}a.\gamma=&0\qquad\qquad\ \ \qquad\qquad\ \mbox{ on }\partial\Omega_{\tau,\tau+2}\\ \phi_{\tau}a(\tau,x)=&0\qquad\qquad\qquad\ \qquad\ \ \ \mbox{ in }\Omega\end{aligned}\right.$
Then from integrability estimation(appendix-11) we have:
$\|\phi_{\tau}a\|_{L^{\infty}(\Omega_{\tau,\tau+2})}\leq$ $C_{IE}(\Omega,N,p)\|\phi^{\prime}_{\tau}a+\phi_{\tau}c\|_{L^{p}(\Omega_{\tau,\tau+1})}$
(31) $\|a\|_{L^{\infty}(\Omega_{\tau+1,\tau+2})}\leq$ $C_{IE}(\Omega,N,p)(1+M_{\phi})K_{c}(1+\tau)^{\mu_{c}}$
Furthermore from the relation $partial_{t}b\leq c$ using jensen inequality we get for some $\mu_{1}\in\mathbb{N}$
$\|b\|_{L^{p}(\Omega)}\leq K_{1}(1+t)^{\mu_{1}}$
Integrating with respect to time variable from $\tau$ to $\tau+1$, we will get some $\mu_{2}\in\mathbb{N}$
$\|b\|_{L^{p}(\Omega_{\tau,\tau+1})}\leq K_{2}(1+\tau)^{\mu_{2}}$
This along with relation-(31) implies for some $\mu_{3}\mathbb{N}$
$\|ab\|_{L^{p}(\Omega_{\tau,\tau+1})}\leq K_{3}(1+\tau)^{\mu_{3}};\qquad p>\frac{N+2}{2}$

### v2

@@ 2. The case of $d_{b}=0$
$\leq$ $\|a_{0}+c_{0}\|_{\mathrm{L}^{p}(\Omega)}\|\phi(0)\|_{\mathrm{L}^{q}(\Omega)}+\|a\|_{L^{p}(\Omega_{0,2})}\|\partial_{t}\phi\|_{L^{q}(\Omega_{0,2})}+d_{a}\|a\|_{\mathrm{L}^{p}(\Omega_{0,2})}\|\Delta\phi\|_{\mathrm{L}^{q}(\Omega_{0,2})}.$
Choose $q=\frac{4.5}{4.5-1}$, i.e., the Hölder conjugate of $4.5$, duality estimate yields
$\|c\|_{\mathrm{L}^{4.5}(\Omega_{0,2})}\leq\|a_{0}+c_{0}\|_{\mathrm{L}^{4.5}(\Omega)}+2C_{q,d_{c}}\|a\|_{\mathrm{L}^{4.5}(\Omega_{0,2})}.$
Choose $K_{0,1}:=\max\left\{\|a_{0}\|_{\mathrm{L}^{\infty}(\Omega)}2^{\frac{2}{9}}|\Omega|^{\frac{2}{9}}+K_{2}K_{7}+\|a_{0}+c_{0}\|_{L^{4.5}(\Omega)}\right\}$, we have the following estimate
$\|a\|_{\mathrm{L}^{4.5}(\Omega_{0,1})}+\|c\|_{\mathrm{L}^{4.5}(\Omega_{0,1})}\leq K_{0,1}.$
@@
$\partial_{t}\left(\phi_{\tau}a\right)-d_{a}\Delta\left(\phi_{\tau}a\right)&=a\partial_{t}\phi_{\tau}+\phi_{\tau}(c-ab)\qquad\ \ \mbox{ in }\Omega_{\tau,\tau+2}\\ \phi_{\tau}\nabla a\cdot n&=0\qquad\qquad\qquad\qquad\qquad\ \mbox{ on }\partial\Omega_{\tau,\tau+2}\\ \phi_{\tau}(\tau)a(\tau,x)&=0\qquad\qquad\qquad\qquad\qquad\ \ \mbox{ in }\Omega.\end{aligned}\right.$
@@
$\partial_{t}\zeta-d_{a}\Delta\zeta&=a\partial_{t}\phi_{\tau}+\phi_{\tau}c\qquad\qquad\mbox{in}\ \Omega_{\tau,\tau+2}\\ \nabla\zeta\cdot n&=0\qquad\qquad\qquad\qquad\mbox{on}\ \partial\Omega_{\tau,\tau+2}\\ \zeta(\tau,x)&=0\qquad\qquad\qquad\qquad\ \ \mbox{in}\ \Omega.\end{aligned}\right.$
@@
Employing the integrability estimate for $\zeta$ available from [2, Lemma 3.3] (see Theorem 11 from the Appendix for the precise statement), we deduce that
@@
$\partial_{t}\left(\phi_{\tau}c\right)-d_{c}\Delta\left(\phi_{\tau}c\right)=&c\partial_{t}\phi_{\tau}+\phi_{\tau}(ab-c)\qquad\mbox{ in}\Omega_{\tau,\tau+2}\\ \phi_{\tau}\nabla c\cdot n&=0\qquad\qquad\qquad\qquad\ \mbox{ on }\partial\Omega_{\tau,\tau+2}\\ \phi_{\tau}(\tau)c(\tau,x)&=0\qquad\qquad\qquad\qquad\ \mbox{ in }\Omega.\end{aligned}\right.$
where $\phi_{\tau}(s):=\phi(s-\tau)$ for $s\in[\tau,\infty)$ with $\tau>0$. Thanks to the positivity of $c$, employing the maximum principle for the heat equation and the integrability estimate from [2, Lemma 3.3] (see Theorem 11 from the Appendix for the precise statement), we arrive at
@@
$\left\|b(t,\cdot)\right\|^{\frac{9}{4}}_{\mathrm{L}^{\frac{9}{4}}(\Omega)}$ $\leq 2^{\frac{5}{4}}(1+t)^{\frac{5}{4}}\left(\left\|b_{0}\right\|_{\mathrm{L}^{\infty}(\Omega)}^{\frac{9}{4}}\,\left|\Omega\right|+(1+t)^{\frac{47}{20}}K_{12}^{\frac{9}{4}}\right)\leq K_{13}(1+t)^{\frac{18}{5}},$
@@
where we have used the bounds obtained earlier. As $\frac{7}{2}<\frac{(5)\frac{3}{2}}{5-3}$, employing the maximum principle for the heat equation and the integrability estimate from [2, Lemma 3.3] (see Theorem 11 for the precise statement) as before, we get
@@
In our above proof Gagliardo-Nirenberg inequality played an important role. We will use the same technique to obtain $\mathrm{L}^{\frac{3}{2}}(\Omega)$ integral estimation of $b$ for dimension $N=1,2$.
@@
Let $(a,b,c)$ be the solution of the degenerate system (2) in dimensions $N=1,2$. Then, there exists positive constants $\hat{K}$ and $K_{c}>0$ and some $\mu_{c}\in\mathbb{N}$, independent of time, such that
@@
The proof is similar to the proof of proposition 3. We use Gagliardo-Nirenberg inequality which says there exists a positive constant depending on the domain and dimension $C_{2}=C_{2}(\Omega,N)$ such that
$\|D^{j}\sqrt{a}\|_{\mathrm{L}^{p}(\Omega)}\leq C_{2}\|D^{m}\sqrt{a}\|_{\mathrm{L}^{r}(\Omega)}^{\alpha}\|\sqrt{a}\|_{\mathrm{L}^{q}(\Omega)}^{1-\alpha}+\tilde{C_{2}}(\Omega,N,r,q)\|\sqrt{a}\|_{L^{q}(\Omega)},$
where
$1\leq p,q,r\leq+\infty\ \mbox{and}\ \frac{1}{p}=\frac{j}{N}+\left(\frac{1}{r}-\frac{m}{N}\right)\alpha\ +\frac{1-\alpha}{q}\ \ \mbox{with}\ \ \frac{j}{N}\leq\alpha\leq 1$
For dimension $N=2$, we choose $r=2,m=1,j=0,q=2$ and $\alpha=\frac{1}{2}$. This particular choice yields the following
$\|\sqrt{a}\|_{\mathrm{L}^{4}(\Omega)}^{4}\leq 2^{3}C_{2}^{4}\|\nabla\sqrt{a}\|_{\mathrm{L}^{2}(\Omega)}^{2}\|\sqrt{a}\|_{\mathrm{L}^{2}(\Omega)}^{2}+2^{3}\tilde{C_{2}}^{4}\|\sqrt{a}\|_{\mathrm{L}^{2}(\Omega)}^{4}.$
Just like in proposition 3, the following estimates hold
$\int_{\tau}^{\tau+2}\|\sqrt{a}\|_{W^{1,2}(\Omega)}^{2}\leq K_{1},\ \ \int_{\tau}^{\tau+2}\|\sqrt{c}\|_{W^{1,2}(\Omega)}^{2}\leq K_{1},\quad\forall\tau\geq 0,$
where $K_{1}=\frac{2E(a_{0},b_{0},c_{0})}{d_{a}}+\frac{2E(a_{0},b_{0},c_{0})}{d_{c}}+2M_{1}|\Omega|+2M_{2}|\Omega|$. We consider the relation we obtained from Gagliardo-Nirenberg inequality. Integrating that relation from time $\tau$ to $\tau+2$ and thanks to our previous estimates, we conclude the following
$\int_{\tau}^{\tau+2}\|\sqrt{a}\|_{\mathrm{L}^{4}(\Omega)}^{4}\leq 2^{4}C_{2}^{4}M_{1}^{2}|\Omega|^{2}K_{1}+2^{4}\tilde{C_{2}}^{4}M_{1}^{4},$
$\int_{\tau}^{\tau+2}\|\sqrt{c}\|_{\mathrm{L}^{4}(\Omega)}^{4}\leq 2^{4}C_{2}^{4}M_{1}^{2}|\Omega|^{2}K_{1}+2^{4}\tilde{C_{2}}^{4}M_{1}^{4}.$
@@
Let’s choose $K_{2}=(2^{4}C_{2}^{4}M_{1}^{2}|\Omega|^{2}K_{1}+2^{4}\tilde{C_{2}}^{4}M_{1}^{4})^{\frac{1}{2}}$. we arrive at the following estimates $\forall\tau\geq 0$.
$\|a\|_{\mathrm{L}^{2}(\Omega_{\tau,\tau+2})}\leq K_{2},\qquad\|c\|_{\mathrm{L}^{2}(\Omega_{\tau,\tau+2})}\leq K_{2}.$
Applying Hölder inequality we arrive at the following estimates
(33) $\|a\|_{\mathrm{L}^{\frac{5}{3}}(\Omega_{\tau,\tau+2})}\leq 2K_{2}|\Omega|^{\frac{1}{6}},\qquad\|c\|_{\mathrm{L}^{\frac{5}{3}}(\Omega_{\tau,\tau+2})}\leq 2K_{2}|\Omega|^{\frac{1}{6}}\quad\forall\tau\geq 0.$
Rest of the proof is similar to the proof of proposition 3. We obtain a time independent constant $\hat{K}>0$ such that for all $t\geq 0$
$\|b\|_{\mathrm{L}^{\frac{3}{2}}(\Omega)}\leq\hat{K}(1+t)^{\frac{5}{6}}.$
For dimension $N=1$, choose $r=2,m=1,j=0,q=2$ and $\alpha=\frac{1}{3}$. Arguing similarly like in the case of dimension $N=2$, we obtain time independent constants $\hat{K}>0,K_{c}>0$ and $\mu_{c}\in\mathbb{N}$ such that for all $\tau\geq 0$
$\|b\|_{\mathrm{L}^{\frac{3}{2}}(\Omega)}$ $\leq\hat{K}(1+\tau)^{\frac{5}{6}}\qquad\ \tau\geq 0,$
$\|a\|_{\mathrm{L}^{3.5}(\Omega_{\tau,\tau+1})}$ $\leq K_{c}(1+\tau)^{\mu_{c}}\qquad\tau\geq 0,$
$\|c\|_{\mathrm{L}^{3.5}(\Omega_{\tau,\tau+1})}$ $\leq K_{c}(1+\tau)^{\mu_{c}}\qquad\tau\geq 0.$
@@
Above Proposition 1 and Proposition 3 will help us to relate entropy dissipation with the missing term $\delta_{B}$. Furthermore the integral estimates of $c$ will provide us polynomial growth of the solutions too. We begin by the polynomial growth of the solutions in the next theorem.
@@
Let $(a,b,c)$ be the solution of degenerate system (2). Furthermore, let for $p>\frac{N+2}{2}$, there exists a time independent constant $K_{c}>0$ and $\mu_{c}\in\mathbb{N}$, such that
$\|a\|_{\mathrm{L}^{p}(\Omega_{\tau,\tau+1})}+\|c\|_{\mathrm{L}^{p}(\Omega_{\tau,\tau+1})}\leq K_{c}(1+\tau)^{\mu_{c}}\quad\forall\tau\geq 0$
Then there exists constant $K_{\infty}$ and $\mu\in\mathbb{N}$, such that
@@
The species $a$ is a positive subsolution of the following equation in the time interval $(0,2)$.
@@
(34) $\partial_{t}a(t,x)-d_{a}\Delta a(t,x)\leq&c\qquad\qquad\qquad\mbox{ in }\Omega_{0,2}\\ \nabla a(t,x)\cdot n=&0\qquad\qquad\quad\ \ \mbox{ on }\partial\Omega_{0,2}\\ a(0,x)=&a_{0}\qquad\quad\qquad\ \ \mbox{ in }\Omega.\end{aligned}\right.$
The solution corresponding to the equation (34) can be expressed as:
$a\leq\int_{\Omega}G_{d_{a}}(t,0,x,y)a_{0}(y)\rm{d}y+\int_{0}^{t}\int_{\Omega}G_{d_{a}}(t,s,x,y)c(s,y)\rm{d}y\rm{d}s.$
where $G_{d_{a}}$ denotes the Green’s function associated with the operator $\partial_{t}-d_{a}\Delta$ with Neumann boundary condition. We use the fact that ${a_{0}(y)\in\mathrm{L}^{\infty}(\Omega)}$ and ${\int_{\Omega}G_{d_{a}}(t,0,x,y)\rm{d}y\leq 1}$ for all $t\in(0,2)$. It yields
$\left|\int_{\Omega}G_{d_{a}}(t,0,x,y)a_{0}(y)\rm{d}y\right|\leq\|a_{0}\|_{\mathrm{L}^{\infty}(\Omega)}.$
We use the following Green’s function estimate is from [10][3]: there exists a constant $K_{1}>0$, depending only on the domain, such that
$0\leq G_{d_{a}}(t_{1},s,x,y)\leq\frac{K_{1}}{(4\pi(t_{1}-s))^{\frac{N}{2}}}e^{-\kappa\frac{\|x-y\|^{2}}{(t_{1}-s)}}=:g_{d_{a}}(t_{1}-s,x-y),$
$\left\|g_{d}\right\|_{\mathrm{L}^{z}(\Omega_{0,2})}\leq K_{2}\qquad\forall z\in\left[1,1+\frac{N}{2}\right)$
for some constant positive $K_{2}$ depending on $z$. We in particular choose ${\frac{1}{z}=1-\frac{1}{p}}$ which is admissible because ${p>\frac{N+2}{2}}$. Hence
$\left|\int_{0}^{t}\int_{\Omega}G_{d_{a}}(t,s,x,y)c(s,y)\rm{d}y\rm{d}s\right|\leq 23^{\mu_{c}}K_{2}K_{c}.$
It further yields the pointwise bound for the species $a$ in the unit parabolic cylinder in initial time. More precisely, we have
(35) $\|a\|_{\mathrm{L}^{\infty}(\Omega_{0,2})}\leq\|a_{0}\|_{\mathrm{L}^{\infty}(\Omega)}+23^{\mu_{c}}K_{2}K_{c}.$
Next we consider the differential relation $\partial_{t}b\leq c$. It yields
(36) $b\leq b_{0}+\int_{0}^{t}c.$
Employing Minkowski’s integral inequality on the above estimate we obtain
$\|b\|_{\mathrm{L}^{p}(\Omega_{0,2})}\leq 2^{\frac{1}{p}}\left(\|b_{0}\|_{\mathrm{L}^{\infty}(\Omega)}|\Omega|^{\frac{1}{p}}+3^{\mu_{c}}2^{\frac{2p-1}{p}}K_{c}\right).$
Using the estimate (35), we derive the following estimate
$\|ab\|_{\mathrm{L}^{p}(\Omega_{0,2})}\leq 2^{\frac{1}{p}}\left(\|a_{0}\|_{\mathrm{L}^{\infty}(\Omega)}+2K_{2}K_{c}\right)\left(\|b_{0}\|_{\mathrm{L}^{\infty}(\Omega)}|\Omega|^{\frac{1}{p}}+3^{\mu_{c}}2^{\frac{2p-1}{p}}K_{c}\right)=:K_{b,[0,1]}.$
Similar to the arguments for the species $a$, the above estimates yields the following pointwise estimate for the species $c$
$\|c\|_{\mathrm{L}^{\infty}(\Omega_{0,2})}\leq\|c_{0}\|_{\mathrm{L}^{\infty}(\Omega)}+3^{\mu_{c}}K_{2}K_{b,[0,1]}=:K_{c,[0,1]}.$
The pointwise estimate of the species $b$ in unit time is a consequence of the above estimate. The relation (36) and the non-negativity of the species $b$, yields
$\|b\|_{\mathrm{L}^{\infty}(\Omega_{0,2})}\leq\|b_{0}\|_{\mathrm{L}^{\infty}(\Omega)}+2K_{c,[0,1]}.$
Consider the following constant ${K_{\infty,[0,2]}:=\max\left\{\|a_{0}\|_{\mathrm{L}^{\infty}(\Omega)}+23^{\mu_{c}}K_{2}K_{c},K_{c,[0,1]},\|b_{0}\|_{\mathrm{L}^{\infty}(\Omega)}+2K_{c,[0,1]}\right\}}$. This helps us deduce
$\|a\|_{\mathrm{L}^{\infty}(\Omega_{0,2})},\|b\|_{\mathrm{L}^{\infty}(\Omega_{0,2})},\|c\|_{\mathrm{L}^{\infty}(\Omega_{0,2})}\leq K_{\infty,[0,2]}.$
Let $\phi:[0,\infty)\rightarrow[0,1]$ be a smooth function such that $\phi(0)=0$, $\phi\big{|}_{[1,\infty)}=1\text{\ and \ }\phi^{\prime}\in[0,M_{\phi}]$ for some positive constant $M_{\phi}$. Let’s denote $\phi_{\tau}(s)=\phi(s-\tau)$. The product function $\phi_{\tau}a$ satisfies the following initial-boundary value problem:
$\partial_{t}\phi_{\tau}a-d_{a}\Delta\phi_{\tau}a=&\phi^{\prime}_{\tau}a+\phi_{\tau}(c-ab)\qquad\mbox{ in }\Omega_{\tau,\tau+2}\\ \nabla\phi_{\tau}a.n=&0\qquad\qquad\ \ \qquad\qquad\ \mbox{ on }\partial\Omega_{\tau,\tau+2}\\ \phi_{\tau}a(\tau,x)=&0\qquad\qquad\qquad\ \qquad\ \ \mbox{ in }\Omega.\end{aligned}\right.$
Applying integrability estimation (see Theorem 11 from the Appendix) on the above equation, we obtain
$\|\phi_{\tau}a\|_{\mathrm{L}^{\infty}(\Omega_{\tau,\tau+2})}\leq$ $C_{IE}(\Omega,N,p)\|\phi^{\prime}_{\tau}a+\phi_{\tau}c\|_{\mathrm{L}^{p}(\Omega_{\tau,\tau+1})}$
(37) $\|a\|_{\mathrm{L}^{\infty}(\Omega_{\tau+1,\tau+2})}\leq$ $C_{IE}(\Omega,N,p)(1+M_{\phi})K_{c}(1+\tau)^{\mu_{c}}.$
Next we consider the differential relation $\partial_{t}b\leq c$. It yields
$b\leq b_{0}+\int_{0}^{t}c.$
Employing Minkowski’s integral inequality on the above estimate we obtain
$\|b\|_{\mathrm{L}^{p}(\Omega_{0,\tau})}\leq(1+\tau)^{\frac{1}{p}}\left(\|b_{0}\|_{\mathrm{L}^{\infty}(\Omega)}|\Omega|^{\frac{1}{p}}+(1+\tau)^{1+\mu_{c}}K_{c}\right).$
We derive the following integral estimate of $ab$
$\|ab\|_{\mathrm{L}^{p}(\Omega_{\tau,\tau+1})}\leq C_{IE}(\Omega,N,p)(1+M_{\phi})K_{c}(1+\tau)^{\mu_{c}}+(1+\tau)^{\frac{1}{p}}\left(\|b_{0}\|_{\mathrm{L}^{\infty}(\Omega)}|\Omega|^{\frac{1}{p}}+(1+\tau)^{1+\mu_{c}}K_{c}\right)$
where $p>\frac{N+2}{2}$. Let us denote the following positive constant ${K_{3}:=C_{IE}(\Omega,N,p)(1+M_{\phi})K_{c}+\left(\|b_{0}\|_{\mathrm{L}^{\infty}(\Omega)}|\Omega|^{\frac{1}{p}}+K_{c}\right)}$. It yields
$\|ab\|_{\mathrm{L}^{p}(\Omega_{\tau,\tau+1})}\leq K_{3}(1+t)^{2+\mu_{c}}\qquad\mbox{where}\ p>\frac{N+2}{2}.$